OpenStudy (anonymous):

Determine the solution of the system. y = x + 6 y = 2x

hero (hero):

Hint: y = y x + 6 = 2x

sam (.sam.):

y = x + 6 y = 2x -------------------- x + 6= 2x -x+6=0 -x=-6 x=-6*-1 x=6 ---------------------- y = 2(6)=12

OpenStudy (anonymous):

A farmer plans to create a rectangular garden that he will enclose with chicken wire. The garden can be no more than 30 ft wide. The farmer would like to use at most 180 ft. of chicken wire. Write a system of linear inequalities that models this situation.

sam (.sam.):

Post a new problem :)

OpenStudy (anonymous):

i just did

OpenStudy (anonymous):

A farmer plans to create a rectangular garden that he will enclose with chicken wire. The garden can be no more than 30 ft wide. The farmer would like to use at most 180 ft. of chicken wire. Write a system of linear inequalities that models this situation.

sam (.sam.):

I mean post it using 'new question', see attachment below

OpenStudy (anonymous):

k